Problem 14.73 17 of 18 〉 Review I Constants Periodic Table Part A An infants toy has a 120 g wooden animal hanging from a spring. If pulled down gently, the animal oscillates up and down with a period of 0.47 S. His older sister pulls the spring a bit more than intended. She pulls the animal 32 cm below its equilibrium position, then lets go. The animal flies upward and detaches from the spring right at the animals equilibrium position. If the animal does not hit anything on the way up, how far above its equilibrium position will it go? Express your answer with the appropriate units. | Value Units Submit Request Answer
Answer #1

Time, T = 2 pi sqrt (m/k)

k = 4 pi^2 m/T^2

k = (4 pi^2 x 0.120)/0.47^2 = 21.45 N/m

Using law of conservation of energy,

0.5 kx^2 = mgh

h = kx^2/2 mg

h = (21.45 x 0.32^2)/(2 x 0.12 x 9.8) = 0.93 m

Comment in case any doubt please rate my answer .. .